# references.bib

### v1

@@ 4.2.1 Cost Analysis
The cost analysis evaluates the total financial expenditure required to process the dataset using different models. The results indicate that ORI offers balanced cost-efficiency, with a total cost of \$0.025997, which is slightly higher than Meta-Llama-3-70B (\$0.023300) but lower than Deepseek-LLM-67B (\$0.033286) and Calme-2.4 (\$0.030512).
@@
Meta-Llama-3-70B 0.023300
Qwen2-72B 0.024964
Deepseek-LLM-67B 0.033286
Calme-2.4 0.030512
ORI 0.025997
Table 5: Total cost comparison across models using the MMLU dataset. ORI demonstrates competitive cost-efficiency relative to other models.
@@ 4.2.2 Token Generation Speed
Token generation speed reflects how quickly each model produces outputs. ORI achieved a speed of 606.93 tokens/second, outperforming Qwen2-72B (480.3 tokens/second) and Calme-2.4 (450 tokens/second), while slightly surpassing Deepseek-LLM-67B (600.0 tokens/second). Although Meta-Llama-3-70B leads with 680.8 tokens/second, the gap is marginal when considering ORI’s superior cost-efficiency and competitive latency.
@@
Meta-Llama-3-70B 680.8
Qwen2-72B 480.3
Deepseek-LLM-67B 600.0
Calme-2.4 450.0
ORI 606.93
Table 6: Token generation speed comparison across models using the MMLU dataset. ORI balances speed and cost efficiency, outperforming several models.
@@ 4.2.3 Latency Analysis
Latency measures the total time taken to process the dataset, reflecting the model’s ability to deliver timely responses. ORI achieved a total latency of 624.01 seconds, which is almost identical to Deepseek-LLM-67B (624.11 seconds) and significantly better than Qwen2-72B (651.84 seconds) and Calme-2.4 (693.45 seconds). Although Meta-Llama-3-70B recorded the lowest latency at 596.37 seconds, the difference is minimal compared to ORI, especially considering ORI’s superior cost and speed performance.
The efficient latency of ORI is a result of its dynamic routing system, which directs tasks to the most suitable models, minimizing bottlenecks. Although Meta-Llama-3-70B offers slightly faster processing, ORI’s balanced performance across all key metrics makes it a more reliable and cost-effective solution for varied tasks.
@@
Meta-Llama-3-70B 596.37
Qwen2-72B 651.84
Deepseek-LLM-67B 624.11
Calme-2.4 693.45
ORI 624.01

### v2

@@ 4.2.1 Cost Analysis
The cost analysis evaluates the total financial expenditure required to process the dataset using different models. The results indicate that ORI offers balanced cost-efficiency, with a total cost of \$1.94, which is slightly higher than Meta-Llama-3-70B (\$1.49) but lower than Deepseek-LLM-67B (\$7.06) and Calme-2.4 (\$2.24).
@@
Meta-Llama-3-70B 1.49
Qwen2-72B 2.35
Deepseek-LLM-67B 7.06
Calme-2.4 2.24
ORI 1.94
Table 5: Total cost comparison across models using the MMLU dataset. ORI demonstrates competitive cost-efficiency relative to other models. [^2] Model and pricing information provided by https://aimlapi.com; the models were accessed via their API for use in the routing model. Only one provider was used for the experiment to ensure an unbiased analysis.
@@ 4.2.2 Token Generation Speed
Token generation speed reflects how quickly each model produces outputs. ORI achieved a speed of 375 tokens/second, outperforming Qwen2-72B (131 tokens/second) and Calme-2.4 (128 tokens/second), while surpassing Deepseek-LLM-67B (60 tokens/second). Although Meta-Llama-3-70B leads with 444 tokens/second, the gap is marginal when considering ORI’s superior cost-efficiency and competitive latency.
@@
Meta-Llama-3-70B 444
Qwen2-72B 131
Deepseek-LLM-67B 60
Calme-2.4 128
ORI 375
Table 6: Token generation speed comparison across models using the MMLU dataset. ORI balances speed and cost efficiency, outperforming several models. [^4] Token generation speeds were calculated by dividing the total number of tokens generated by the total processing time (in seconds).
@@ 4.2.3 Latency Analysis
Latency measures the total time taken to process the dataset, reflecting the ability of the model to deliver timely responses. ORI achieved a total latency of 562 seconds, which is significantly better than Qwen2-72B (605 seconds) and Calme-2.4 (608 seconds). The efficient latency of ORI is a result of its dynamic routing system, which directs tasks to the most suitable models, minimizing bottlenecks. Although Meta-Llama-3-70B offers slightly faster processing, ORI’s balanced performance across all key metrics makes it a more reliable and cost-effective solution for varied tasks.
@@
Meta-Llama-3-70B 531
Qwen2-72B 605
Deepseek-LLM-67B 731
Calme-2.4 608
ORI 562